mike and dan wanted to buy a present for their parents anniversary The anniversary was only one month away
every day i will put one of my coins in this piggy bank and you put one of yours in it toomike said ok said dan." ihave 22 nickels in my wallet" i have 18 dimes in mine" said mike
one day the brothers were surprised to notice dan actually had more money than mike in his wallet. how many days had passed since they had begun putting their coins into the piggy bank?

Answer by scott8148(5891)

let d=days

to find the day that they have equal amounts ___ 5(22-d) = 10(18-d)

distributing ___ 110 - 5d = 180 - 10d

adding 10x and subtracting 110 ___ 5d = 70

dividing by 5 ___ d = 14

after the 14th day, Dan will have more than Mike